The vertices of a triangle are represented by the complex numbers 4 - 2i, -1 + 4i, and 6 + i, then the complex number representing the centroid of a triangle is
A. 3 + i
B. 3 - i
C. 9 + i
D. 9 - i

Grade:Upto college level

1 Answers

Sunil Raikwar
9 years ago
centroid of the triangle is (z1+z2+z3)/3=(4-2i-1+4i+6+i)/3= 3+i
Thanks & Regards
Sunil Raikwar
askIITians Faculty